John wants to travel from P to B in India. He had two options to reach
B from P. He can either travel by air which takes 30 minutes or by a taxi
which takes 10 hours. The air ticket will cost him Nu. 5,000 and the taxi fare will costs him Nu. 1,250. Other things remaining constant, John would gain by choosing air travel if, and only if, his time were valued at more than:

Answer by stanbon(75887)

John wants to travel from P to B in India. He had two options to reach
B from P. He can either travel by air which takes 30 minutes or by a taxi
which takes 10 hours.
The air ticket will cost him Nu. 5,000 and the taxi fare will costs him Nu. 1,250.
------
Unit cost by air: Nu 5000/(1/2) = Nu 10000 per hour
==================
Unit cost by taxi: Nu 1250/10 = Nu 125 per hour
=======================================================


Other things remaining constant, John would gain by choosing air travel if, and only if, his time were valued at more than (NU 10,000-Nu 125) = Nu 875
====================
Cheers,
Stan H.
====================